Please share in detail


1.     SOSU, Inc has before tax income this year is $900,000. The company’s payout ratio is 40%. The company's common equity currently has a book value of $5,000,000. They just paid a dividend of $1.87, and the required rate of return on this stock is 10%. Compute the value of this stock if dividends are expected to continue growing indefinitely at the company's internal growth rate. Tax rate = 28%.

 

 


1
Expert's answer
2021-04-19T18:46:07-0400

We will find it by the following formula:

"P=\frac{EPS(1-RR)}{r-(RR\times ROE)}"

"DPR=\frac{D}{net income}"

net income (EPS)=4.675

RR=1-DPR=1-0.4=0.6

"ROE=\frac{900 000-0.28\times900 000}{5 000 000}=0.1296"

"P=\frac{4.675(1-0.6)}{0.1-(0.6\times 0.1296)}=84.08"